OpenStudy (tia_sweetheart17):

factor completely. 5x^2+53x+72

OpenStudy (raden):

hint : 5 = 5 * 1 72 = 8 * 9 5 * 9 + 1 * 8 = 53 nah, continue it ..

OpenStudy (anonymous):

So first we can establish that this is a quadratic equation which can be solved multiple ways: - X-method -Quadratic formula -Completing the square My method of chose is the quadratic formula just for the simplicity of things.. \[(-b \pm \sqrt{b^2-4ac})/2a\] Which you then plug in your numbers.. And will get two answers, you will then check both answers to make sure they are correct.